# for Academic Information Seeking

### v1

@@ Abstract
Applying large language models (LLMs) for academic API usage shows promise in reducing researchers’ academic information seeking efforts. However, current LLM API-using methods struggle with complex API coupling commonly encountered in academic queries. To address this, we introduce SoAy, a solution-based LLM API-using methodology for academic information seeking. It uses code with a solution as the reasoning method, where a solution is a pre-constructed API calling sequence. The addition of the solution reduces the difficulty for the model to understand the complex relationships between APIs. Code improves the efficiency of reasoning.
@@
To evaluate SoAy, we introduce SoAyBench, an evaluation benchmark accompanied by SoAyEval, built upon a cloned environment of APIs from AMiner. Experimental results demonstrate a 34.58-75.99% performance improvement compared to state-of-the-art LLM API-based baselines. All datasets, codes, tuned models, and deployed online services are publicly accessible at https://github.com/RUCKBReasoning/SoAy.
[Uncaptioned image] A Solution-based LLM API-using Methodology
for Academic Information Seeking
Yuanchun Wang, Jifan Yu§ [^1] Work was done when these authors interned at Zhipu AI., Zijun Yao§, Jing Zhang, Yuyang Xie§, Shangqing Tu§ Yiyang Fu, Youhe Feng, Jinkai Zhang, Jingyao Zhang, Bowen Huang, Yuanyao Li Huihui Yuan, Lei Hou§, Juanzi Li§and Jie Tang§ Renmin University of China      §Tsinghua University      Zhipu AI https://soay.aminer.cn
1 Introduction
@@
Researchers frequently seek academic information such as a paper’s metadata, a scholar’s publication record, and the connections between scholars as part of their routine scholarly activities. The advent of academic information retrieval systems—including DBLP dblp team (2005), CiteSeer Giles et al. (1998), and Google Scholar Noruzi (2005)—and more sophisticated academic mining platforms such as Microsoft Academic Search Sinha et al. (2015) and AMiner Tang et al. (2008), has substantially reduced the complexity of exploring the academic data, which encompasses massive entities like authors, conferences, and papers.
@@
Figure 1: Comparing responses from Chameleon Lu et al. (2023), ToolLLaMA Qin et al. (2023b), and the proposed SoAy, all leveraging AMiner APIs to address the same academic query.
Large Language Models Zhao et al. (2023) (LLMs) unveil the potential to further reduce the effort required for academic information seeking, benefiting from LLMs’ reasoning ability to comprehend intentions expressed in natural language. Enabling LLMs to utilize existing academic retrieval/mining APIs Qin et al. (2023a) tends to be a feasible solution. However, this adaptation is not easy because of two main challenges:
(1) Coupling: Academic information for academic queries is distributed across various storage locations but interconnected internally, resulting in multiple academic APIs responsible for different functions yet interconnected, i.e. coupled. This requires LLM to fully understand the sequential order of calls between these APIs. Existing API-using methods like Chameleon Lu et al. (2023), shown in Figure 1(a), retrieve and execute APIs separately to get results, thus failing to handle such API coupling.
(2) Efficiency: Existing methods such as ToolLLM Qin et al. (2023b) illustrated in Figure 1(b) employ a depth-first search-based decision tree (DFSDT) for step-by-step reasoning, which has the potential to capture the regularities in the sequence of API. However, the excessive steps lead to query times that users may find unacceptable, compared with traditional academic search engines that offer researchers a very fast response.
We propose SoAy, a Solution-based API-using methodology to apply LLMs to for academic information seeking. Our key innovation lies in the new API calling method: solution and code. As shown in Figure 1(c), SoAy first lets the LLM generate a feasible API calling plan, i.e. solution, based on complex user inputs, and then allows the LLM to generate executable API calling code based on the generated solution. The pre-generated solutions can assist the LLM in clarifying the complex coupling relationships between academic APIs when generating subsequent code. Calling APIs through code generation in a single inference process without repeated reasoning can achieve efficiency levels closer to traditional academic search systems. Specifically, the ability for LLM to reliably utilize complex academic APIs in this manner to accomplish academic search tasks benefits from the automatic generation of a training dataset, consisting of (query, solution, code) triplets. The data is created on the AMiner APIs, resulting in 3,960 automated triplets. We retain 4/5 of these triplets, providing valuable resources for fine-tuning existing LLMs or demonstrating LLM in-context learning. This alignment helps bridge the gap between general LLMs and the intricacies of academic API usage.
To evaluate SoAy, we utilize 1/5 of the generated data, comprising 792 (query, solution, code) triplets. These are enhanced through human verification to ensure the quality of the test set. We then introduce SoAyEval, an evaluation methodology that takes into account both solution correctness and code execution accuracy. By employing SoAyEval for automated evaluation alongside human assessment, our experiments demonstrate the effectiveness and efficiency of SoAy in utilizing AMiner APIs compared to state-of-the-art LLM API-using baselines, across both in-context learning and tuning modes. Furthermore, we validate the versatility of SoAy for application in other academic retrieval systems.
The paper contributes: (1) SoAy, a novel LLM adaptation method, addressing challenges like coupling and efficiency in academic information seeking through automatic data generation, teaching LLMs to use academic APIs. (2) SoAyBench, a released benchmark dataset, accompanied by the SoAyEval evaluation method, to assess LLMs’ proficiency in using academic APIs. (3) Demonstrated effectiveness and efficiency of SoAy in the AMiner system, along with its real deployment on an online service as shown in Figure 13.
Figure 2: Illustration of AMiner API Library, API Graph and solution library.
2 Task Formulation
The task for SoAy is to seek appropriate academic information $I$ based on complicated user query $Q$ in an efficient way. This objective tends to be achieved by using academic APIs in an valid and efficient order.
Problem 1
User Intent Parse Based on $Q$, SoAy needs to understand which academic information the user requires, then identify the associations within them and conduct separate queries.
$Q\rightarrow Q_{S},\quad Q_{S}=\{Q_{i}\}_{i=1}^{N}.$ (1)
Traditional academic search systems cannot accomplish this task, thus requiring users to break their requirements $Q$ down into step-by-step queries $Q_{S}$ to type in the search box. SoAy, on the other hand, needs to use LLMs to break the intent down and get the relevant API calling sequence.
Problem 2
Academic Information Aggregation After understanding what information the user truly needs, SoAy’s task is to retrieve relevant information from data sources according to $Q_{S}$ and aggregate and process the results to provide the user with the precise information $I$ they seek.
Traditional information retrieval systems lack this design and instead display all retrieved information to users for them to select what they want in a list-like format. What SoAy needs to do is generate the correct code to call the academic APIs and process the obtained information within a comparable time to those traditional systems.
@@ 2 Task Formulation
Figure 3: The overall process of SoAy around one case of an academic question. The main steps of data generation are shown on the left. Two mode of LLM alignment with the automated triplet dataset are on the right.
3 SoAy: Solution-based API-using
@@
We introduce SoAy to adapt general LLMs for academic information seeking, emphasizing the use of existing academic APIs. To achieve this, abundant training data in the form of (academic query, executable API sequence) pairs is essential, facilitating either fine-tuning of existing LLMs or guiding them through in-context learning. Illustrated in the left part of Figure 3, this data generation pipeline comprises four key steps:
@@
(1) Solution Library Construction. Utilizing the provided academic APIs, we construct a graph representing the dependencies among APIs, and traverse this graph to identify valid and simple paths, referred to as solutions. These solutions collectively form a library known as the solution library.
@@
(2) Combination Formulation. To obtain a more precise semantic of each solution in the library, we combine all the possible inputs of the initial API and all the outputs of the final API of the solution to formulate a series of combinations.
@@
(3) Query Generation. We employ ChatGPT to transform each combination into a template question lacking specific academic entity data. These templates are then diversified into various expressions and populated into potential academic queries with data from a specialized academic source, the AMiner database here.
(4) Code Generation. Finally, we use ChatGPT again to write a piece of corresponding code given the combination, which is expected to correspond to the template query that shares the same combination. Lastly, we employ ChatGPT once more to generate the corresponding code for the given combination. This code is anticipated to match the template query that has the same combination.
Notably, in the data preparation part, we indicate ChatGPT as gpt-3.5-turbo-0613. We present the details of the data preparation prompt and the process of setting up SoAyBench in the appendix.
4 Solution Library Construction
@@
$f_{n}$ $\xrightarrow{a_{n,m}}f_{m},$ (2)
$if\quad a_{n,m}=i_{m}$ $\cap o_{n}\quad\text{and}\quad a\neq\emptyset,$
where $i_{n}$ denotes the input parameters of $f_{m}$ and $o_{n}$ denotes the returned attributes of $f_{n}$.
[…]
Definition 2
A Solution $S$ is defined as an execution sequence of APIs traversed from the API graph:
@@
Valid Solution. Among all traversed solutions from the API graph, validity is determined by starting with a fuzzy API, as only fuzzy APIs can accept user utterances in natural language as input.
Simple Solution. For a single query, there may be multiple solutions available. The preferred choice is the simplest one, containing the fewest APIs. For instance, when inquiring about the publication date of the representative work of a scholar, the optimal solution is “searchPerson $\rightarrow$ getPersonPubs”, despite obtaining the same information is also possible through “searchPerson $\rightarrow$ getPersonPubs $\rightarrow$ getPublication $\rightarrow$ searchPublication”.
@@
We meticulously craft prompts to instruct ChatGPT in fulfilling the above two characteristics. All prompts for question generation are detailed in Appendix Part III. Prompts. Consequently, we’ve created 16 solutions to establish the solution library $L_{S}$ for AMiner APIs. Figure 2(c) presents a subset of these solutions, including their input parameters, returned attributes, and corresponding question templates.
5 Reasoning Code Generation
In order for LLM to aggregate effective academic information and provide users with precise answers, SoAy let LLM generate code containing API calls for academic information retrieval. To achieve this, a large amount of Query, Corresponding Code data needs to be generated to train LLM.
5.1 Combination Formulation
@@
Section 2 describes the one-to-several relationship between a query and multiple solutions. It’s evident that a similar one-to-several relationship exists between a solution and several query intents. In a solution $S$, the first API is referred to as the head API $f_{1}$, while the last API is termed the tail API $f_{J}$. For example, in Figure 2(c), consider the 3-hop solution “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo”. Due to the diversity of the Head API Input $i_{1}$ and the Tail API Output $o_{J}$, determining which information (outputs) is sought under what conditions (inputs) remains uncertain even after the solution is established. An intuitive approach is to combine the head API input, the solution, and the tail API output to refine the intent space into a specific requirement. A solution-based combination is defined as a specific binding of the head API input, the solution, and the tail API output: $C=(i_{1},S,o_{J})$. Given the $i_{1}$ is “title” and the $o_{J}$ is “education_experience”, the $S$ “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo” could lead to one plausible interpretation: querying the graduation school of the first author of the provided publication. It’s important to recognize that the information from the academic data source corresponding to $i_{1}$ is both the information needed by q-t to fill and the input information needed by the generated code, $o_{J}$ is both the expected answer of q-t and the golden result of the code.
5.2 Query Generation
Template-Query Generation. For a given combination $C$, we define a template question $Q_{T}$ to represent the intent of $C$ with missing entity information. In the example provided, the corresponding template question $Q_{T}$ could be formulated as “Which school did the first author of XXXX graduate from?”, where “XXXX” refers to a placeholder representing the title of a publication, as illustrated in Table 1.
Semantic Expansion and Entity Filling. A specific intent represented by a template question $Q_{T}$ may be expressed by multiple natural language queries. For instance, “Which school did XXX graduate from?” and “What is XXX’s alma mater?” convey the same intent. To enrich the semantic diversity of an intention, we conduct semantic augmentation by crafting prompts to leverage ChatGPT in generating three new, semantically enriched template questions for each combination.

### v2

# SoAy: A Solution-based LLM API-using Methodology for Academic Information Seeking

@@ Abstract.
Applying large language models (LLMs) to academic API usage shows promise in reducing researchers’ efforts to seek academic information. However, current LLM methods for using APIs struggle with the complex API coupling commonly encountered in academic queries. To address this, we introduce SoAy, a solution-based LLM methodology for academic information seeking. SoAy enables LLMs to generate code for invoking APIs, guided by a pre-constructed API calling sequence referred to as a solution. This solution simplifies the model’s understanding of complex API relationships, while the generated code enhances reasoning efficiency. LLMs are aligned with this solution-oriented, code-based reasoning method by automatically enumerating valid API coupling sequences and transforming them into queries and executable code.
@@
To evaluate SoAy, we introduce SoAyBench, an evaluation benchmark accompanied by SoAyEval, built upon a cloned environment of APIs from AMiner. Experimental results demonstrate a 34.58-75.99% performance improvement compared to state-of-the-art LLM API-based baselines. All codes are publicly accessible at https://github.com/RUCKBReasoning/SoAy.
Large Language Model, API, Academic Information Seeking
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709412isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Multi-agent planningccs: Computing methodologies Semantic networksccs: Information systems Data mining
1. Introduction
@@
Researchers frequently seek academic information such as a paper’s metadata, a scholar’s publication record, and the connections between scholars as part of their routine scholarly activities. The advent of academic information retrieval systems—including DBLP (dblp team, 2005), CiteSeer (Giles et al., 1998), and Google Scholar (Noruzi, 2005)—and more sophisticated academic mining platforms such as Microsoft Academic Search (Sinha et al., 2015) and AMiner (Tang et al., 2008), has substantially reduced the complexity of exploring the academic data, which encompasses massive entities like authors, conferences, and papers.
However, existing systems are constrained by their predefined functionalities and may not fully address the flexible and complex needs of users. For example, when a researcher aims to know how many times the representative work of Yann LeCun has been cited, he or she often handcrafts the search keywords into Google Scholar search box, finds and clicks the target item on the result page, and finally seeks the target publication and its citation count. Large Language Models (LLMs) (Zhao et al., 2023), with their robust reasoning capabilities, offer a promising solution for comprehensively understanding and addressing such nuanced requirements. One potential approach involves enabling LLMs to leverage academic seeking APIs (Qin et al., 2023a). The diverse functionalities offered by these APIs, when combined effectively, hold the potential to address a wide array of user needs. However, this adaptation presents two significant challenges:
(1) Coupling: One key challenge in adapting large language models (LLMs) to leverage academic retrieval/mining APIs is the inherent parameter coupling within these APIs. For instance, to find the most cited paper of a scholar, the process involves identifying the scholar, retrieving their papers, and then sorting by citation count. This sequence requires the LLM to understand the interdependent calls between the scholar and publication search APIs (i.e., searchPerson and getPersonPubs APIs). Existing methods, such as Chameleon (Lu et al., 2023), as depicted in Figure 1(a), retrieve and execute APIs independently, overlooking the necessary coupling. Consequently, they fail to correctly invoke the getPersonPubs API, which requires a person ID provided by the searchPerson API.
(2) Efficiency: Existing methods, such as ToolLLM (Qin et al., 2023b), employ a depth-first search-based decision tree (DFSDT) for step-by-step reasoning, which can effectively capture the couplings in the sequence of API calls. However, the numerous steps involved result in multiple inference times, which may be unacceptable to users who are accustomed to the rapid response of existing academic information retrieval systems. For instance, as shown in Figure 1(b), to obtain the most cited paper of a scholar, DFSDT requires one LLM inference per returned paper from the getPersonPubs API to determine its citation count, leading to high inference costs.
@@
Figure 1. Comparing responses from Chameleon (Lu et al., 2023), ToolLLaMA (Qin et al., 2023b), and the proposed SoAy, all leveraging AMiner APIs to address the same academic query.
[…]
Comparing responses from Chameleon (Lu et al., 2023), ToolLLaMA (Qin et al., 2023b), and the proposed SoAy, all leveraging AMiner APIs to address the same academic query.
@@
Figure 2. The overall SoAy process involves a single academic question case. The main steps of data generation are illustrated on the left and middle sections. Two modes of LLM alignment with the automated triplet dataset are on the right.
[…]
The overall SoAy process involves a single academic question case. The main steps of data generation are illustrated on the left and middle sections. Two modes of LLM alignment with the automated triplet dataset are on the right.
We propose SoAy, a Solution-based API-using methodology to apply LLMs for academic information seeking. Our key innovation is the solution-oriented, code-based method for API calling. As shown in Figure 1(c), SoAy first lets the LLM generate a feasible API calling plan, i.e. solution, based on complex user inputs, and then allows the LLM to generate executable API calling code based on this solution. The pre-generated solutions help the LLM clarify complex coupling relationships between academic APIs when generating subsequent code. By generating the API calling code in a single inference process, SoAy achieves efficient inference without repeated reasoning. To enable LLMs to reliably generate solutions and code in this manner, it is crucial to automatically create a training dataset consisting of (Query, Solution, Code) triplets. The process involves using academic APIs published by aminer.org to create a solution library by constructing a dependency graph among the APIs and identifying valid API calling paths. Each solution is then linked with input parameters and output attributes to align more precisely with user intentions. This binding is transformed into a template query and then into Python code. Code correctness is ensured by instantiating the input, executing the code, and verifying the results against expected outputs.
As a result, we automatically develop a dataset containing 3,960 (query, solution, code) triplets, consisting of SoAyBench. We then introduce SoAyEval, an evaluation methodology that takes into account both solution correctness and code execution accuracy. The experiments demonstrate the effectiveness and efficiency of SoAy in utilizing AMiner APIs compared to state-of-the-art LLM API-using baselines, across both in-context learning and tuning modes. Furthermore, we validate the versatility of SoAy for application in other academic retrieval systems.
The paper contributes: (1) SoAy, A novel LLM adaptation method that addresses challenges such as coupling and efficiency in academic information seeking. This method involves automatic data generation, enumerating API calling sequences, transforming them into queries and executable code, and teaching LLMs to use academic APIs. (2) SoAyBench, a released benchmark dataset, accompanied by the SoAyEval evaluation method, to assess LLMs’ proficiency in using academic APIs. (3) Demonstrated effectiveness and efficiency of SoAy in the AMiner system, along with its real deployment on an online service.
2. SoAy: Solution-based API-using
2.1. Background
The process users seek academic information using traditional academic information retrieval systems can be summarized as follow:
Intent Decomposition: To address their complex information needs, users often begin by formulating keywords and entering them into a search interface. For example, to seek the answer of the question illustrated in Figure 1, a user might input “New York University Yann Lecun” into the search box of the academic information retrieval system to explore relevant results.
Step-by-step Action: Users then follow a step-by-step process to find the information they need. This involves searching in the search box, browsing through the retrieved scholars, selecting the target scholar, and exploring further by continued browsing and clicking.
This sequence of user actions essentially corresponds to a series of API calls that underpin traditional academic systems. For instance, entering a query in the search box translates to calling the searchPerson API as depicted in Figure 3. Clicking on a scholar in the search results corresponds to calling APIs like getPersonPubs to retrieve more detailed scholar information. Finally, clicking on a paper with the most citations to find out its citation count involves calling the getPublication API. Thus, in this scenario, the API calling sequence is: searchPerson $\rightarrow$ getPersonPubs $\rightarrow$ getPublication. This sequence of API calls is defined as a solution, as detailed in Definition 3.2.
Figure 3. Illustration of AMiner API Library $L_{A}$, API Graph $G$ and solution library $L_{S}$.
[…]
Illustration of AMiner API Library $L_{A}$, API Graph $G$ and solution library $L_{S}$.
2.2. SoAy Framework
@@
To tailor general LLMs to comprehend user requirements and methodically invoke academic APIs correctly and efficiently to access the desired information, we utilize pre-collected solutions $S$ to aid the LLM in crafting Python code that incorporates the necessary API calls. To achieve this, abundant training data in the form of (Query, Solution, Code) triplet is essential, facilitating either fine-tuning of existing LLMs or guiding them through in-context learning. Illustrated in the left and middle parts of Figure 2, this solution collection and data generation pipeline comprises four key steps:
@@
(1) Solution Library Construction (Section 3). Utilizing the provided academic APIs, we construct a graph representing the dependencies among APIs, and traverse this graph to identify valid and concise paths, referred to as solutions. These solutions collectively form a library known as the solution library.
@@
(2) Combination Formulation (Section 4.1). To obtain more precise semantics of each solution in the library, we combine all the possible inputs of the initial API and all the outputs of the final API of the solution to formulate a series of combinations.
@@
(3) Query Generation (Section 4.2). We employ ChatGPT to transform each combination into a template question lacking specific academic entity data. These templates are then diversified into various expressions and transformed into specific academic queries, populated with concrete academic entities as input and output from the specialized academic source, the AMiner database.
(4) Code Generation (Section 4.3). Finally, we use ChatGPT to generate Python code corresponding to the given combination, which is expected to handle various queries that share the same combination. For each instantiated query, we ensure code correctness by incorporating the query input into the code, executing the code, and verifying the results against the query’s ground truth outputs.
Notably, during the process of data creation, we indicate ChatGPT as gpt-3.5-turbo-0613.
3. Solution Library Construction
@@
(1) $f_{n}$ $\xrightarrow{a_{n,m}}f_{m},$
$if\quad a_{n,m}=i_{m}$ $\cap o_{n}\quad\text{and}\quad a_{n,m}\neq\emptyset,$
where $i_{m}$ denotes the input parameters of $f_{m}$ and $o_{n}$ denotes the returned attributes of $f_{n}$.
The API coupling $f_{n}\xrightarrow{a_{n,m}}f_{m}$ means that we can use $f_{n}$ to output the attribute $a_{n,m}$, and subsequently use $a_{n,m}$ as the input parameter to call $f_{m}$. This is because $a_{n,m}$ is the intersection of the returned attributes $o_{n}$ of $f_{n}$ and the input parameters $i_{m}$ of $f_{m}$. The graph $G$ built from AMiner APIs according to their coupling relationships is shown in Figure 3(b). By traversing this graph while limiting the maximum number of hops to $H$, we can derive an API calling sequence, referred to as a solution:
Definition 0.
A Solution $S$ is defined as a calling sequence of APIs traversed from the API graph:
@@
Valid Solution. A valid solution can be checked by judging whether the node in the API graph corresponding to the first API in this solution has zero Indegree. Note this is specific to academic information seeking systems. The situation may vary for other application systems.
Concise Solution. For a single query, there may be multiple solutions available. The preferred choice is the most concise one, containing the fewest APIs. We utilize ChatGPT to verify if two given solutions can obtain the same information. The specific prompt used for this verification can be found in our code repo. Specifically, each traversed solution is sequentially added to the solution library $L_{S}$. For each new solution $S$, we check whether a more concise solution already exists in $L_{S}$ that can achieve the same result as $S$.
@@
Consequently, we’ve created 16 solutions to establish the solution library $L_{S}$ for AMiner APIs. Figure 3(c) presents a subset of these solutions, including their input parameters, returned attributes, and corresponding question templates.
4. Soluion-based Data Creation
For each solution in the solution library, $L_{S}$, we generate corresponding queries and code. This enables LLMs to effectively perform academic information seeking.
4.1. Combination Formulation
@@
Typically, a solution may correspond to multiple query intents, each representing a distinct need. Therefore, we formulate combinations derived from a solution to clearly represent each specific intent. We here define a head API $f_{1}$ as the first API in a solution $S$, while the last API is termed the tail API $f_{J}$. For example, in Figure 3(c), considering the 3-hop solution “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo”, searchPublication is $f_{1}$ and $f_{J}$ is getPersonBasicInfo. Due to multiple choices of the Head API Input $i_{1}$ and the Tail API Output $o_{J}$, determining which information (outputs) is sought under what conditions (inputs) remains uncertain even after the solution is established. We define a solution-based combination as a specific binding of the head API input, the solution, and the tail API output: $C=(i_{1},S,o_{J})$, to refine the solution $S$ into a targeted intent. For example, given $i_{1}$ is “title” and $o_{J}$ is “education_experience”, the solution “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo” can refer to a specific interpretation: querying the alma mater of the first author of the provided publication.
4.2. Query Generation
Template-Query Generation. For each combination $C$, we define a template question $Q_{T}$ to capture its intent. For instance, for the aforementioned combination C=C=(“title”, “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo”, “education_experience”), the template question $Q_{T}$ could be phrased as, “Which school did the first author of XXXX graduate from?” where “XXXX” serves as a placeholder for the “title of a publication”, as shown in Table 1. Additionally, $Q_{T}$ could be expressed in variations like, “What is the alma mater of the first author of XXXX?” to convey the same intent. To broaden the semantic diversity of an intention, we engage in semantic augmentation by prompting ChatGPT to generate three varied template questions for each combination, enhancing the range of phrasing and nuances.